OpenStudy (bryanbarrett02):

What is the value of the expression? 3 -5 4 • 4 A. 4−15 B. 4−2 C. 42 D. 48

Nnesha (nnesha):

\[4^3 \cdot 4^{-5}\] use `^` to represent exponents exponent rule: \[\large\rm x^n \cdot x^m = x^{n+m}\] when we multiply same base we should add their exponents.

OpenStudy (bryanbarrett02):

ok

Nnesha (nnesha):

ok let me know what you get :=))

OpenStudy (bryanbarrett02):

64*1024

Nnesha (nnesha):

uh huh don't use calculator.

OpenStudy (bryanbarrett02):

why not

Nnesha (nnesha):

just add the exponents here is an example \[\huge\rm 2^4 \cdot 2^3 = 2^{4+3}=2^7\]

OpenStudy (bryanbarrett02):

? but thats not a b c or d

Nnesha (nnesha):

because thats what they want u to do

Nnesha (nnesha):

what do you mean ??

OpenStudy (bryanbarrett02):

u said 2^7

Nnesha (nnesha):

in this expression `4^3` base is 4 and exponent is 3 `4^{-5}` base is 4 and exponent is -5 because bases are the same and there is multiplication sign you have to add the exponents

Nnesha (nnesha):

yes because 4+3 =7 correct ?

OpenStudy (bryanbarrett02):

i get that but that cant be the answer if its not one of the option

Nnesha (nnesha):

i have mentioned that is an example ...

Nnesha (nnesha):

do the same thing \[4^3 \cdot 4^{-5}\] for this question

OpenStudy (bryanbarrett02):

ok i feel dum now lol

Nnesha (nnesha):

no you're not. :=))

Nnesha (nnesha):

alright add the exponents what would you get ?

OpenStudy (bryanbarrett02):